Bubba’s Western Wear is a western hat retailer in Lubbock, Texas. Although Bubba’s carries numerous styles of western hats, each hat has approximately the same price and invoice (purchase) cost, as shown in the following table. Sales personnel receive a commission to encourage them to be more aggressive in their sales efforts. Currently, the Lubbock economy is really humming, and sales growth at Bubba’s has been great. The business is very competitive, however, and Bubba, the owner, has relied on his knowledgeable and courteous staff to attract and retain customers who otherwise might go to other western wear stores. Because of the rapid growth in sales, Bubba is also finding the management of certain aspects of the business more difficult, such as restocking inventory and hiring and training new salespeople. Sales price $ 80.00 Per unit variable expenses Purchase cost 23.50 Sales commissions 31.50 Total per unit variable costs $ 55.00 Total annual fixed

expenses Advertising $ 297,500 Rent 248,000 Salaries 397,000 Total fixed expenses $ 942,500 Required: 1. Calculate the annual breakeven point, both in terms of units and in terms of sales dollars. 2. If Bubba’s sells 39,000 hats, what is its before-tax income or loss? Support your answer by preparing a contribution income statement. 3. If Bubba’s sells 52,000 hats, what is its margin of safety (MOS) in dollars and MOS ratio? 4. Bubba is considering the elimination of sales commissions completely and increasing salaries by $357,000 annually. What wo
Answer #1

Given Data:

Selling Price $                      80 per hat
Total Variable cost $                      55 per hat
Contribution $                      25 per hat
Fixed Cost $            942,500 per year

1) Annual break even Point (BEP):

BEP = Total Fixed Cost /Contribution per unit

BEP in terms of units = $942,000/25 = 37,700 units

BEP in terms of Dollar = Total units sold x Selling price per unit

BEP in terms of Dollar = 37,700 x 80 = $ 3,016,000

2) If Bubba's sell 39000 hats

Total Income = Total Units sold x Selling price per unit

i.e. =39000 x $80 = $ 3,120,000

Total Variable cost = Total Units sold x variable cost per unit

i.e. = 39000 x $55 = $2,145,000

Total Fixed cost = $942,500 (given in question)

INCOME STATEMENT

INCOME $ EXPENSES $

Sales 3,120,000 Variable cost 2,145,000

Fixed Cost 942,500

Profit before tax 32,500   

Total 3,120,000 Total 31,20,000

If Bubba's sale 39,000 units then he will earned a profit before tax $ 32,500

3) Margin of Safety (MOS)

Current Sales                  52,000 hats
BEP sales                  37,700 hats
Total Current sales $        4,160,000 (52,000 x $80)
Break Even sales $        3,016,000 (37,700 x $80)

MOS = ((Current Sales Level-Break even sales Level /Current Sales level ) x 100

i.e. ($4,160,000-$3,016,000/$4,160,000) x 100 = 27.50%

MOS in Dollars = $4,160,000-$3,016,000 = $1,144,000

MOS in Units = 52,000 - 37,700 = 14,300 hats

4) If, Bubba is considering the elimination of sales commissions completely and increasing salaries by $357,000 annually then BEP would be:

Revised data will be in given question

Selling Price $          80.00 per hat
Total Variable cost $          23.50 per hat - No sales commission
Contribution $          56.50 per hat
Fixed Cost $ 1,299,500 per year - including salary of Rs. 3,57,000

BEP = Total Fixed Cost /Contribution per unit

BEP in terms of units = $1,299,500/56.50 = 23,000 units

BEP in terms of Dollar = Total units sold x Selling price per unit

BEP in terms of Dollar = 23,000 x $ 80 = $ 1,840,000